jagr2713 (jagr2713):

Trick Question: Maria have some blueberries and some cups. If she put 9 blueberries into each cup she will have two cups left. If she put 6 blueberries into each cup she will have three blueberries left over. Explain

jagr2713 (jagr2713):

@pooja195 @rational

OpenStudy (anonymous):

because its divided by 3?

OpenStudy (rational):

are we finding number of blueberries and cups ?

pooja195 (pooja195):

B(blueberries) = 9(C - 2) C=cups B(blueberries) = 6C + 3 c= cups C = 7, B = 45. There are 7 cups and 45 blueberries this is messy ;-;

pooja195 (pooja195):

btw @Nnesha thanks for teaching me these types of prolems before :)

jagr2713 (jagr2713):

finding the leftover of berries

jagr2713 (jagr2713):

Easier way: The number of blueberries is equal to number of cups multiplied by 4 minus one left. That is : B = 9(C - 2) -> (1) Then the number of blueberries will be equal to the number of cups times three plus one left. That is : B = 6C + 3 -> (2) Now substitute the 1st and 2nd equations: B = 9(C - 2) B = 6C + 3 C = 7, B = 45. Therefore Blueberries = 45 and cups = 7.

OpenStudy (rational):

Neat :)

OpenStudy (campbell_st):

don't over hand the blueberries... then it would be messy

jagr2713 (jagr2713):

lol

pooja195 (pooja195):

LOL

pooja195 (pooja195):

jinx xD